OpenStudy (anonymous):

Solve the equation -3(y-5)=15 y = ?

jimthompson5910 (jim_thompson5910):

I'll get you started -3(y-5)=15 -3(y-5)/(-3) = 15/(-3) ... Divide both sides by -3. y - 5 = -5 y = ???

OpenStudy (anonymous):

@jim_thompson5910 0??

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

@jim_thompson5910 thank you so much!!!

jimthompson5910 (jim_thompson5910):

Check: -3(y-5)=15 -3(0-5)=15 ... plug in y = 0 -3(-5) = 15 15 = 15 ... works

jimthompson5910 (jim_thompson5910):

yw